# Ground States of ${\text{Class}\;\mathcal{S}}$ Theory on ADE Singularities and dual Chern-Simons theory

### v1

@@ B.2 Contribution from half-hypermultiplet $I^{trifund}$
where the prefactor 2 comes from 2 choices of the first SU$(2)$ trifundamental index.
Case 2: $\mu=1,\nu=-1$. The first constraint implies
$n_{1}-n_{2}=0\mod{4},$
and the second constraint says that summing over all such pairs will help us take care of the case $\mu=-1,\nu=1$ case. The Boltzmann weight sum is
$F_{1,-1}=2\frac{1+t^{12}}{(1-t^{6})(1-t^{12})}.$
Adding these up, we have
$I^{trifund}_{155}=(F_{11}+F_{1,-1})(t^{2}-t^{4})=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (83)
In the case of $I^{trifund}_{255}$, the constraints are
$\sum e^{\pi i/2(n_{1}-n_{2})}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu}$ $=\sum e^{\pi(\mu+\nu)i/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu},$
$\sum e^{\pi in_{2}}c^{\mu\nu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{\mu\nu}$ $=-\sum(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-\mu,-\nu}.$
@@
The only difference from the previous (155) case is the minus sign in the second equation, but this does not affect the number of independent solutions. This suggests that
$I^{trifund}_{255}=I^{trifund}_{155}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (84)
In the case of $I^{trifund}_{355}$, the constraints are
$\sum e^{\pi i/2(n_{1}-n_{2})}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu}$ $=-\sum e^{\pi(\mu+\nu)i/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu},$
$\sum e^{\pi in_{2}}c^{\mu\nu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{\mu\nu}$ $=\sum(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-\mu,-\nu},$
from which we have
$e^{\pi i/2(n_{1}-n_{2})}c^{\mu\nu n_{1}n_{2}}$ $=-e^{\pi(\mu+\nu)i/2}c^{\mu\nu n_{1}n_{2}},$
$e^{\pi in_{1}}c^{-\mu,-\nu,n_{2}n_{1}}$ $=(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}.$
Case 1: $\mu=\nu=1$. The first constraint implies
$n_{1}-n_{2}=0\mod{4}.$
Case 2: $\mu=1,\nu=-1$. The first constraint implies
$n_{1}-n_{2}=2\mod{4}.$
Notice that these are the same constraints as the (155) case, except the order is different. Therefore,
$I^{trifund}_{355}=I^{trifund}_{155}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (85)
@@
$I^{trifund}_{455}=I^{trifund}_{155}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (86)
In summary, we have the result that, for the nonabelian holonomies that appear twice in the triplet,
$I^{trifund}_{155}=I^{trifund}_{255}=I^{trifund}_{355}=I^{trifund}_{455}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (87)
@@
$\sum e^{\pi i/2(n_{1}-n_{2})}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu\rho}$ $=\sum e^{\pi i/2(\mu+\nu+\rho)}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu\rho},$
$\sum e^{\pi in_{2}}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{\mu\nu\rho}$ $=\sum(-1)^{(\mu+\nu+\rho-3)/2}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-\mu,-\nu,-\rho},$
from which we have
$e^{\pi i/2(n_{1}-n_{2})}c^{\mu\nu\rho n_{1}n_{2}}$ $=e^{\pi i/2(\mu+\nu+\rho)}c^{\mu\nu\rho n_{1}n_{2}},$
$e^{\pi in_{1}}c^{-\mu,-\nu,-\rho n_{2}n_{1}}$ $=(-1)^{(\mu+\nu+\rho-3)/2}c^{\mu\nu\rho n_{1}n_{2}}.$
Case 1: $\mu=1,\nu=1,\rho=1$. In this case we have
$n_{1}-n_{2}=3\mod{4},$
and the sum yields
$F_{+++}=\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}.$
Case 2: $\mu=1,\nu=1,\rho=-1$. In this case we have
$n_{1}-n_{2}=1\mod{4},$
and the sum yields
$F_{++-}=\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}=F_{+++}.$
Case 3: $\mu=1,\nu=-1,\rho=1$. In this case we have
$n_{1}-n_{2}=1\mod{4},$
and the sum yields
$F_{+-+}=\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}=F_{+++}.$
Case 4: $\mu=1,\nu=-1,\rho=-1$. In this case we have
$n_{1}-n_{2}=3\mod{4},$
and the sum yields
$F_{+--}=\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}=F_{+++}.$
@@
$I^{trifund}_{555}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (88)
Now, we have all results we need to compute the supersymmetric Casimir energy.
@@ B.3 Supersymmetric Casimir Energy computation
When we sum over all the indices to compute the Casimir energy, we need to multiply the vector multiplet contribution by a factor of $1/2$, since each vector multiplet is shared between two Trinions. For example, for the triplet (111), we have
$I_{111}$ $=I^{trifund}_{111}+\frac{3}{2}I^{vec}_{1}$
$=-\frac{1}{18\beta}+\frac{9}{2}-\frac{35\beta}{54}+O(\beta^{2}),$
which suggests that the supersymmetric Casimir energy is
@@
$I_{112}$ $=I^{trifund}_{112}+\frac{3}{2}I^{vec}_{1}$
$=-\frac{1}{18\beta}+\frac{9}{2}-\frac{251\beta}{54}+O(\beta^{2}),$ (91)
$I_{114}$ $=I_{113}=I^{trifund}_{113}+\frac{3}{2}I^{vec}_{1}$
$=-\frac{1}{18\beta}+\frac{9}{2}-\frac{251\beta}{54}+O(\beta^{2})$ (92)
@@
$I_{115}$ $=I^{trifund}_{115}+I^{vec}_{1}+\frac{1}{2}I^{vec}_{5}$
$=-\frac{1}{18\beta}+3-\frac{143\beta}{54}+O(\beta^{2}),$ (95)
@@
$I_{155}$ $=I^{trifund}_{155}+\frac{1}{2}I^{vec}_{1}+I^{vec}_{5}$
$=-\frac{1}{18\beta}+\frac{3}{2}-\frac{35\beta}{54}+O(\beta^{2}).$ (98)
@@
There is a single state: (555), that is left to consider.
$I_{555}$ $=I^{trifund}_{555}+\frac{3}{2}I^{vec}_{5}$
$=-\frac{1}{18\beta}+\frac{19}{54}\beta+O(\beta^{2}).$ (101)

### v2

@@ B.2 The contribution $I^{\text{(trifund)}}$ of the half-hypermultiplet
where the prefactor 2 comes from 2 choices of the first SU$(2)$ trifundamental gauge index, and $t^{2}-t^{4}$ is the contribution of the fields $q^{1}$ and $\bar{\psi}_{\dot{+}}$.
In the case of $I^{\text{(trifund)}}_{255}$, the constraints are
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu}$ $=\sum e^{i\pi(\mu+\nu)/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu},$
$\sum(-1)^{n_{2}}c^{\mu\nu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{\mu\nu}$ $=-\sum(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-\mu,-\nu}.$
@@
The only difference from the previous (155) case is the minus sign in the second equation, but this does not affect the number of independent solutions. Thus,
$I^{\text{(trifund)}}_{255}=I^{\text{(trifund)}}_{155}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (95)
In the case of $I^{\text{(trifund)}}_{355}$, the constraints are
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu}$ $=-\sum e^{i\pi(\mu+\nu)/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu},$
$\sum(-1)^{n_{2}}c^{\mu\nu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{\mu\nu}$ $=\sum(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-\mu,-\nu},$
from which we obtain
$e^{i\pi(n_{1}-n_{2})/2}c^{\mu\nu n_{1}n_{2}}=-e^{i\pi(\mu+\nu)/2}c^{\mu\nu n_{1}n_{2}},\qquad(-1)^{n_{1}}c^{-\mu,-\nu,n_{2}n_{1}}=(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}.$ (96)
  • If $\mu=\nu=1$, the first constraint implies
    $n_{1}-n_{2}\equiv 0\pmod{4}.$
  • If $\mu=1,\nu=-1$, the first constraint implies
    $n_{1}-n_{2}\equiv 2\pmod{4}.$
Noting that these are for the same constraints as the (155) case [equations (92) and (93)] except that (92) goes with $\nu=-1$ this time and (93) goes with $\nu=1$, we immediately get
$I^{\text{(trifund)}}_{355}=I^{\text{(trifund)}}_{155}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (97)
@@
$I^{\text{(trifund)}}_{455}=I^{\text{(trifund)}}_{155}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (98)
In summary, for the nonabelian holonomies that appear twice in the triplet, we find
$I^{\text{(trifund)}}_{155}=I^{\text{(trifund)}}_{255}=I^{\text{(trifund)}}_{355}=I^{\text{(trifund)}}_{455}=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (99)
@@
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu\rho}$ $=\sum e^{i\pi(\mu+\nu+\rho)/2}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu\rho},$
$\sum(-1)^{n_{2}}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{\mu\nu\rho}$ $=\sum(-1)^{(\mu+\nu+\rho-3)/2}c^{\mu\nu\rho n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-\mu,-\nu,-\rho},$
from which we derive
$e^{i\pi(n_{1}-n_{2})/2}c^{\mu\nu\rho n_{1}n_{2}}$ $=e^{i\pi(\mu+\nu+\rho)/2}c^{\mu\nu\rho n_{1}n_{2}},$
$(-1)^{n_{1}}c^{-\mu,-\nu,-\rho n_{2}n_{1}}$ $=(-1)^{(\mu+\nu+\rho-3)/2}c^{\mu\nu\rho n_{1}n_{2}}.$
  • For $\mu=1,\nu=1,\rho=1$, we have $n_{1}-n_{2}\equiv 3\pmod{4}.$
  • For $\mu=1,\nu=1,\rho=-1$, we have $n_{1}-n_{2}\equiv 1\pmod{4}$.
  • For $\mu=1,\nu=-1,\rho=1$, we have $n_{1}-n_{2}\equiv 1\pmod{4}.$
  • For $\mu=1,\nu=-1,\rho=-1$, we have $n_{1}-n_{2}\equiv 3\pmod{4}.$
For each of the above four cases, the number of pairs $n_{1},n_{2}$ that satisfy the mod 4 constraint, with given $n_{1}+n_{2}$, is the coefficient of $t^{3(n_{1}+n_{2})}$ in the expansion of
$\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}.$
@@
$I^{\text{(trifund)}}_{555}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (100)
Now, we have all the components we need to compute the supersymmetric Casimir energy.
@@ B.3 Supersymmetric Casimir Energy computation
When we sum over all the indices to compute the Casimir energy, we need to multiply the vector multiplet contribution by a factor of $1/2$, since each vector multiplet is shared between two Trinions. We also set $t=\exp\beta$ and expand near $\beta=0$. For example, for the triplet (111) we have
$I_{111}$ $=I^{\text{(trifund)}}_{111}+\frac{3}{2}I^{\text{(vec)}}_{1}=-\frac{1}{18\beta}+\frac{9}{2}-\frac{35\beta}{54}+O(\beta^{2}),$
which leads, using (23), to the supersymmetric Casimir energy contribution from the (111) vertex,
@@
$I_{112}$ $=I^{\text{(trifund)}}_{112}+\frac{1}{2}\left(2I^{\text{(vec)}}_{1}+I^{\text{(vec)}}_{1}\right)=-\frac{1}{18\beta}+\frac{9}{2}-\frac{251\beta}{54}+O(\beta^{2}),$ (103)
$I_{114}$ $=I_{113}=I^{\text{(trifund)}}_{113}+\frac{1}{2}\left(2I^{\text{(vec)}}_{1}+I^{\text{(vec)}}_{3}\right)=-\frac{1}{18\beta}+\frac{9}{2}-\frac{251\beta}{54}+O(\beta^{2}).$ (104)
@@
$I_{115}$ $=I^{\text{(trifund)}}_{115}+I^{\text{(vec)}}_{1}+\frac{1}{2}I^{\text{(vec)}}_{5}=-\frac{1}{18\beta}+3-\frac{143\beta}{54}+O(\beta^{2}),$ (107)
@@
$I_{155}$ $=I^{\text{(trifund)}}_{155}+\frac{1}{2}I^{\text{(vec)}}_{1}+I^{\text{(vec)}}_{5}=-\frac{1}{18\beta}+\frac{3}{2}-\frac{35\beta}{54}+O(\beta^{2}).$ (110)
@@
There is one more state, (555), that is left to consider.
$I_{555}$ $=I^{\text{(trifund)}}_{555}+\frac{3}{2}I^{\text{(vec)}}_{5}=-\frac{1}{18\beta}+\frac{19}{54}\beta+O(\beta^{2}).$ (113)